# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ C.4.1 Part (i)
where the second inequality follows from the conditions that $\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert=o_{p}(1)$ and (4.3). Next, we show $R_{1}-\mathbb{E}(R_{1})=o_{p}(1)$. Conditioning on ${\cal I}_{k}^{c}={\cal I}\backslash{\cal I}_{k}$, we calculate the mean and variance for $R_{1}-\mathbb{E}(R_{1})$:
@@ C.4.2 Part (ii)
From Lemma C.1 and the rate conditions (4.4a), (4.4b) and (4.4c) in Theorem 4.3, we have
@@ Appendix D Details and Examples of the Calibration Procedure
This section provides details and illustrations for the calibration procedure introduced in Section 5. We start with some remarks about the implementation of our calibration procedure. First, it’s important to have the ratio of the sample sizes between the proxy source and target data be equal to that of the original source and the target data. This can be accomplished by downsampling one of the two proxy data. Relatedly, to make the comparisons fairer, it’s useful to rescale the standard error estimate in the transported CI from the calibration procedure by multiplying it with $\sqrt{|{\cal I}_{s_{2}}|/n_{t}}$ in order to mimic the length of the CI for the original TATE. Algorithm 2 provides a step-by-step procedure for calibrating the sensitivity parameters. Figure D.1 exemplifies the calibration regions for estimating the ad effect in Philadelphia in 2024.
@@
1:Source data and confidence level $1-\alpha$
@@
3:Step 2 (Construct CI via the standard approach: With data in ${\cal I}_{s_{2}}$, estimate the ATE and its $(1-\alpha)$ confidence interval, denoted as $\widehat{\text{CI}}_{s_{2}}(1-\alpha)$.
4:Step 3 (Construct CI via our transfer learning approach) :
5:if $|{\cal I}_{s_{1}}|/|{\cal I}_{s_{2}}|\neq n_{s}/n_{t}$ then
6:     Randomly sample from ${\cal I}_{s_{2}}$ a subsample of size ${\cal I}_{t}\cdot(n_{t}/n_{s})$, which without loss of generality assumed to be an integer. Denote the resulting set of indices as ${\cal I}_{s_{2}}$.
@@
8:With $\{({\bf X}_{i},A_{i},Y_{i},S_{i}=1):i\in{\cal I}_{s_{1}}\}\cup\{({\bf V}_{i},S_{i}=0):i\in{\cal I}_{s_{2}}\}$, estimate the ATE on ${\cal S}_{2}$ and its standard error with any $(\gamma_{0},\gamma_{1})\in{\cal C}_{\text{all}}$, denoted as $\widehat{\theta}_{s_{1}\to s_{2}}(\gamma_{0},\gamma_{1})$ and $\widehat{\text{SE}}_{s_{1}\to s_{2}}$. Denote the re-scaled confidence interval as
@@
9:Step 4 (Find the plausible range) : Find the plausible range of sensitivity parameters when transporting from ${\cal S}_{1}$ to ${\cal S}_{2}$:
@@
10:Step 5 (Calibration in the other direction) Exchange ${\cal S}_{1}$ and ${\cal S}_{2}$ and repeat Steps 1-4, resulting in the plausible range ${\cal C}_{2}$.
11:Intersect two plausible regions to construct the final region: ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$.
@@
Figure D.1: The calibration procedure for the analysis in Philadelphia in 2024. Panels from left to right plots ${\cal C}_{1}$, ${\cal C}_{2}$ and ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$ in shadowed areas along $\exp(\gamma_{1})$ in the y-axis and $\exp(\gamma_{0})$ in the x-axis.
We illustrate the overall analysis pipeline for estimating the ad effect in Philadelphia in 2024. First, estimate the TATE across $(\gamma_{0},\gamma_{1})$’s using approaches introduced in Section 4. Specifically, for each $(\gamma_{0},\gamma_{1})$, one will have a $(1-\alpha)$ CI for the TATE. From the left panel of Figure D.2, the ad effect could be insignificant, significant and positive, or significant and negative for Philadelphia voters in 2024 under different choices of $(\gamma_{0},\gamma_{1})$’s. Second, calibrate the sensitivity parameters using Algorithm 2 (i.e., the calibration procedure introduced in Section 5); see middle panel of Figure D.2. Finally, interpreting the ad effect estimates obtained from the first step within the calibrated region obtained from the second step. From the right panel of Figure D.2, this calibration procedure rules out the possibility of having significant and negative effect in Philadelphia in 2024. Therefore, we say Philadelphia is sensitive for a positive effect only. Figure D.3 provides another example on the analysis pipeline for Fulton.
Figure D.2: Analysis pipeline for estimating the ad effect in Philadelphia in 2024. First, estimate the TATE across sensitivity parameters using approaches in Section 4. Note that we only included a few choices of $(\gamma_{0},\gamma_{1})$’s for illustration purposes. Second, calibrate the range of sensitivity parameters using the procedure in Section 5. Finally, overlay the two planes and interpret the TATE within the valid region. Final results are interpreted in Section 6.
Figure D.3: Analysis pipeline for estimating the ad effect in Fulton in 2024. The figure layout is identical to Figure D.2.
@@ E.1 Additional Data Description
E.2 Values of County-Level Ad Effect with the OR Estimator

### v2

@@ C.5.1 Part (i)
where the second inequality follows from the conditions that $\lVert\widehat{\rho}^{(k)}_{a}({\bf V}_{i})-\rho^{(k)}_{a}({\bf V}_{i})\rVert=o_{p}(1)$ and (6). Next, we show $R_{1}-\mathbb{E}(R_{1})=o_{p}(1)$. Conditioning on ${\cal I}_{k}^{c}={\cal I}\backslash{\cal I}_{k}$, we calculate the mean and variance for $R_{1}-\mathbb{E}(R_{1})$:
@@ C.5.2 Part (ii)
From Lemma C.2 and the rate conditions (7a), (7b) and (7c) in Theorem 4.3, we have
@@ Appendix D Details and Examples of the Calibration Procedure
This section provides details and illustrations for the calibration procedure introduced in Section 5.
D.1 Analysis Pipeline
We start with some remarks about the implementation of our calibration procedure. First, it’s important to have the ratio of the sample sizes between the proxy source and target data be equal to that of the original source and the target data. This can be accomplished by downsampling one of the two proxy data. Relatedly, to make the comparisons fairer, it’s useful to rescale the standard error estimate in the transported CI from the calibration procedure by multiplying it with $\sqrt{|{\cal I}_{s_{2}}|/n_{t}}$ in order to mimic the length of the CI for the original TATE. This was mentioned in Algorithm 1 under Step 1. Third, one should make sure the shared covariates in constructing $\widehat{\text{CI}}_{s\to t}(\Gamma_{0},\Gamma_{1};1-\alpha)$ should match the shared covariates ${\bf V}_{i}$ in the actual target sample. See Algorithm 3 for the implementation and Section D of the Supplementary Materials for more discussions. Algorithm 3 provides a step-by-step procedure for calibrating the sensitivity parameters. As an example, Figure D.1 illustrates ${\cal C}_{1}$, ${\cal C}_{2}$ and the final calibration region ${\cal C}$ for estimating the ad effect in Philadelphia in 2024.
@@ D.1 Analysis Pipeline
1:Source data, confidence level $1-\alpha$, set ${\cal C}_{\text{all}}\in\mathbb{R}\times\mathbb{R}$.
@@
3:if ${\cal I}_{s_{1}}/{\cal I}_{s_{2}}>n_{s}/n_{t}$ then
4:     Randomly subset ${\cal I}_{s_{1}}$ of size $|{\cal I}_{s_{2}}|\cdot n_{s}.n_{t}$ and denote the resulting set of indices as ${\cal I}_{s_{1}}$.
5:else
6:     Randomly subset ${\cal I}_{s_{2}}$ of size $|{\cal I}_{s_{1}}|\cdot n_{t}/n_{s}$ and denote the resulting set of indices as ${\cal I}_{s_{2}}$.
@@
8:Step 2.1 (Construct CI via the standard approach: With data in ${\cal I}_{s_{2}}$, estimate the ATE and its $(1-\alpha)$ confidence interval, denoted as $\widehat{\text{CI}}_{s_{2}}(1-\alpha)$.
9:Step 2.2 (Construct CI via our transfer learning approach) :
10:With $\{({\bf X}_{i},A_{i},Y_{i},S_{i}=1):i\in{\cal I}_{s_{1}}\}\cup\{({\bf V}_{i},S_{i}=0):i\in{\cal I}_{s_{2}}\}$, estimate the ATE on ${\cal S}_{2}$ and its standard error with any $(\Gamma_{0},\Gamma_{1})\in{\cal C}_{\text{all}}$, denoted as $\widehat{\theta}_{s_{1}\to s_{2}}(\Gamma_{0},\Gamma_{1})$ and $\widehat{\text{SE}}_{s_{1}\to s_{2}}$. Denote the re-scaled confidence interval as
@@
11:Step 3 (Find the plausible range) : Find the plausible range of sensitivity parameters when transporting from ${\cal S}_{1}$ to ${\cal S}_{2}$:
@@
12:Calibration in the other direction Exchange ${\cal S}_{1}$ and ${\cal S}_{2}$ and repeat Steps 1-3, resulting in the plausible range ${\cal C}_{2}$.
13:Intersect two plausible regions to construct the final region: ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$.
@@
Figure D.1: The calibration procedure for the analysis in Philadelphia in 2024. Panels from left to right plots ${\cal C}_{1}$, ${\cal C}_{2}$ and ${\cal C}={\cal C}_{1}\cap{\cal C}_{2}$ in shadowed areas along $\Gamma_{1}$ in the y-axis and $\Gamma_{0}$ in the x-axis.
D.2 Interpretations
The sensitivity parameters $\Gamma_{0}$ and $\Gamma_{1}$ quantify the change in turnout from 2020 to 2024 in the control arm and the treatment arm, respectively, and different values of $\Gamma_{0}$ and $\Gamma_{1}$ will generally correspond to different effect sizes and direction. Some examples are listed below and Table D.1 enumerates more examples.
  1. 1.
    Suppose $\Gamma_{0}=1$ and $\Gamma_{1}>1$ (i.e., the $y>1$ part in Figure D.1). Then the turnout in 2024 if voters are not exposed to anti-Trump ads will be the same as that in 2020, but the turnout in 2024 if the voters are exposed to anti-Trump ads will be larger than that in 2020. Also, the ad effect in 2024 will be higher than that in 2020 and if $\Gamma_{1}$ is sufficiently large, the effect will be positive and statistically significant.
  2. 2.
    Suppose $\Gamma_{0}>1$ and $\Gamma_{1}=1$ (i.e., the $x>1$ part in Figure D.1). Then the turnout in 2024 if voters are not exposed to anti-Trump ads will be higher than that in 2020, but the turnout in 2024 if the voters are exposed to anti-Trump ads will be the same as that in 2020. Also, the ad effect in 2024 is likely smaller than that in 2020, and if $\Gamma_{0}$ is large enough, the effect may be negative and significant.
  3. 3.
    Suppose $\Gamma_{0}>1$ and $\Gamma_{1}>1$ (i.e., top right region of Figure D.1). Then, the odds of turnout in both treatment and controls will be higher in 2024 than those in 2020. In this case, the ad effect in 2024 may be similar to that in 2020, especially if the shift in the turnouts between 2024 and 2020 are comparable between the control and treatment arms. A similar phenomena would occur if $\Gamma_{0}<1$ and $\Gamma_{1}<1$ (i.e., bottom left region of Figure D.1).
  4. 4.
    Suppose $\Gamma_{0}<1$ and $\Gamma_{1}>1$ (i.e., top left region of Figure D.1). Then, the odd of turnout if voters are not exposed to anti-Trump ads will be lower in 2024 than that in 2020, but the odd of turnout if voter are exposed to anti-Trump ads will be higher than 2024 than that in 2020. Then, the combined effect of the changes in the odds would be a large and positive value of the ad effect in 2024.
  5. 5.
    Suppose $\Gamma_{0}>1$ and $\Gamma_{1}<1$ (i.e., bottom right region of Figure D.1). Then, the odd of turnout if voters are not exposed to anti-Trump ads will be higher in 2024 than that in 2020, but the odd of turnout if voter are exposed to anti-Trump ads will be lower than 2024 than that in 2020. Then, the combined effect of the changes in the odds would be a negative ad effect in 2024 that is large in magnitude.
Table D.1: Examples on the signs of $\Gamma_{0}$, $\Gamma_{1}$ and the ad effect in 2024 (i.e., TATE) compared with the ad effect in 2020.
$\Gamma_{0}$ $\Gamma_{1}$ Odd of turnout in 2024 if unexposed to negative ads (i.e., $Y^{(0)}$) Odd of turnout in 2024 if exposed to negative ads (i.e., $Y^{(1)}$) ATE in 2024 (i.e., TATE)
0 $>0$ same as the corresponding odd in 2020 higher than the corresponding odd in 2020 higher than ATE in 2020
$>1$ $=1$ the corresponding odd in 2020 same as the corresponding odd in 2020 lower than ATE in 2020
$>1$ $>1$ higher than the corresponding odd in 2020 higher than the corresponding odd in 2020 may be similar with ATE in 2020
$<1$ $>1$ lower than the corresponding odd in 2020 higher than the corresponding odd in 2020 higher than ATE in 2020
$>1$ $<1$ higher than the corresponding odd in 2020 lower than the corresponding odd in 2020 lower than ATE in 2020
As discussed in Section 5, not all values of $\Gamma_{0},\Gamma_{1}$ are meaningful and the calibration procedure, which produces the set $\mathcal{C}$ (i.e., the green area in Figure D.1) allows us to focus on values of $\Gamma_{0}$ and $\Gamma_{1}$ that are more interpretable.
@@ E.1 Additional Data Description
E.2 Details for County-Level Ad Effects in Section 6.2
E.2.1 Numeric Values for Figure 3